OpenStudy (anonymous):

Simplify the expressions. You do not have to solve exponents greater than 3. 7^0 16^1 5^2x5^6= 7^9 divided by 7^7 (2x3)^2 2negative^3

OpenStudy (anonymous):

7^0 = 1 16^1 = 16 5^2x5^6= 5^(2+6) = 5^8 7^9 divided by 7^7 = 7^(9-7) = 9^2=81 (2x3)^2=6^2=36 2negative^3 = 1/(2^3)=1/8 If you need a recap f the basic rules, http://www.mathsisfun.com/algebra/exponent-laws.html

OpenStudy (anonymous):

whoops the 4th one is 7^2 = 49 I mixed up the numbers by accident

OpenStudy (anonymous):